# Searching for periodic signals and quasi-periodic oscillations from an extremely active cycle of FRB 20240114A

### v1

@@ Abstract
radio transit source — radio bursts — time series analysis
facilities: FASTsoftware: numpy (Harris et al., 2020), matplotlib (Hunter, 2007), astropy (Astropy Collaboration et al., 2022), PINT (Luo et al., 2021; Susobhanan et al., 2024), emcee (Foreman-Mackey et al., 2013), riptide (Morello et al., 2020), presto (Ransom, 2011), stingray (Huppenkothen et al., 2019), lmfit (Newville et al., 2014), PyEMD (Laszuk, 2017)
1 Introduction
Fast radio bursts (FRBs), first discovered in 2007 (Lorimer et al., 2007), are millisecond-duration, highly energetic radio transients of extragalactic origin. These transient and energetic bursts exhibit dispersion measures (DMs) that significantly exceed the Galactic contribution, thereby strongly suggesting their cosmological distances and placing them among some of the most intriguing astrophysical phenomena (Lorimer et al., 2007; Thornton et al., 2013; CHIME/FRB Collaboration et al., 2021), though large host-galaxy DM contributions have also been found in some cases (e.g., FRB 20190520B; Niu et al. 2022a). Since their initial discovery, the field of FRB research has grown rapidly, with extensive observational campaigns leading to the identification of hundreds of these mysterious bursts. These detections have revealed a diverse population, with some FRBs appearing to repeat sporadically while others have only been observed as singular, non-repeating events (Petroff et al., 2019; CHIME/FRB Collaboration et al., 2021; Xu et al., 2023). The identification of repeating FRBs, most notably FRB 20121102A (Spitler et al., 2016), has ruled out the early speculations that all FRBs originate from cataclysmic one-time events. Instead, the existence of repeaters indicates that at least some FRBs are powered by non-catastrophic engines.
A breakthrough in the field of FRB research was marked by the detection of FRB 20200428, which was definitively associated with an X-ray burst from the Galactic magnetar SGR J1935+2154 (Bochenek et al., 2020; Mereghetti et al., 2020; Li et al., 2021a). Another targeted multi-wavelength campaign did not detect further coincident radio bursts from this source, suggesting that such FRB-SGR associations are rare and might require extreme physical conditions to produce coherent radio emission (Lin et al., 2020). Extensive observations of repeating FRBs have revealed non-Poissonian burst clustering, long-term variations in DM and rotation measure (RM), as well as complex waiting time statistics, highlighting their highly stochastic nature and posing challenges to single rotating compact object models (Li et al., 2021b; Hu & Huang, 2023; Zhang et al., 2024; Liu et al., 2025). Recent polarization and magnetospheric studies suggest that repeating FRBs reside in complex, evolving environments consistent with young magnetars, as indicated by rapid RM variations and strong circular polarization, although observations of FRB 20220912A, which shows high circular polarization but nearly constant RM, suggest that at least some repeaters may not be embedded in strongly magneto-ionic environments (Xu et al., 2022; Feng et al., 2022; Zhu et al., 2023; Feng et al., 2024, 2025). A successful search for pulsar-like, spin-modulated burst-to-burst temporal periodicities in FRB sources would provide a convincing observational evidence that at least some FRBs originate from compact stars.
Some FRB sources have demonstrated distinct long-timescale periodic activity. Notable examples include FRB 20180916B, which follows a 16-days cycle of activity (Chime/Frb Collaboration, 2020; Pastor-Marazuela et al., 2021; Sand et al., 2023), and FRB 20121102A, which has shown signs of a much longer periodicity of approximately 160 days, alongside a shorter periodicity of around 4.6 days (Rajwade et al., 2020; Cruces et al., 2021; Li et al., 2024; Braga et al., 2025). The recently discovered FRB 20240209A, located at the outskirts of a quiescent elliptical galaxy (z = 0.1384), exhibits a periodic activity cycle of about 126 days, further adding to the rare class of periodically repeating FRBs (Pal, 2025). Furthermore, FRB 20220529 shows a potential periodicity of approximately 200 days in its RM, which is consistent with a binary origin, as suggested by the significant RM variation and its rapid recovery (Liang et al., 2025). Similarly, FRB 20201124A exhibits a periodicity of 26 days in its RM, further supporting the idea that periodic RM variations may be a common feature of actively repeating FRBs in binary systems (Xu et al., 2025). The origin of periodic activity in repeating FRBs remains debated, with proposed explanations falling into several categories. Rotation-powered models suggest the periodicity reflects the spin of isolated magnetars (Xu et al., 2021; Lan et al., 2024). Binary models attribute the modulation to orbital dynamics in neutron star systems, including Be/X-ray binaries (Ioka & Zhang, 2020; Lyutikov et al., 2020; Zhang & Gao, 2020; Li et al., 2021c; Barkov & Popov, 2022). Precession effects have also been proposed as a possible cause of the observed periodicity (Yang & Zou, 2020; Tong et al., 2020; Levin et al., 2020; Zanazzi & Lai, 2020; Katz, 2020; Chen, 2020; Chen et al., 2021; Sridhar et al., 2021; Wei et al., 2022). Alternative explanations involve asteroid belt interactions (Dai & Zhong, 2020; Voisin et al., 2021; Deng et al., 2024) or the repetitive collapse of the crust of an accreting strange quark star (Geng et al., 2021). The diversity of models highlights the need for further observational constraints, as many fundamental questions about the physical origins and emission mechanisms of FRBs remain open (Zhang, 2023).
Short-timescale periodicity in FRBs provides critical constraints on their emission mechanisms and progenitors. Some bursts have shown millisecond-scale quasi-periodic structures, e.g. the 0.411 ms spacing in FRB 20201020A (Pastor-Marazuela et al., 2023), although its significance remains marginal. These sub-second periodicities challenge the spin explanations due to extreme rotational requirements, favoring magnetospheric processes instead. Additionally, the detection of ultra-narrow spike components in FRB 20230607A, down to 0.3 ms, further supports a magnetospheric origin for at least some repeating FRBs (Zhou et al., 2025). On the other hand, other studies have found no significant periodicity in repeating FRBs like FRB 20201124A, ruling out isolated magnetars or binaries with certain parameters (Niu et al., 2022b). Recent claims of a $\sim$1.7 s period in FRB 20201124A (Du et al., 2025), if confirmed, could indicate a pulsar origin, while low-significance periodicity candidates such as the 7.267 ms signal in FRB 20230708A (Dial et al., 2025) highlight the challenges in distinguishing true periodicity from burst microstructure. These controversial results highlight the need for higher-cadence and higher time resolution observations to clarify the nature of short-timescale FRB variability.
@@ 1 Introduction
This study systematically analyzes the periodicity of FRB 20240114A, an extremely active FRB. The source was discovered by the Canadian Hydrogen Intensity Mapping Experiment (CHIME) in January 2024, and since then, it has been observed by multiple telescopes (Shin & CHIME/FRB Collaboration, 2024; Tian et al., 2024; Xie et al., 2024; Huang et al., 2025). The source has been continuously observed by the FRB key project of FAST since January 2024. The analysis is based on a dataset collected from January 28 to August 30, 2024, during which 11,553 burst events were recorded (Zhang Junshuo et al. 2025; Zhang Longxuan et al. 2025; Wang et al. 2025, all in preparation). This extensive sample offers the most comprehensive dataset to date for investigating the presence of multiple timescale periodicities in FRBs, aiding in the identification of periodic signals and QPOs from the source. This paper is organized as follows. In Section 2, we search for periodic signals in the TOAs data of FRB 20240114A to explore possible links to magnetar rotation, binary orbits, and precession scenarios. In Section 3, we mainly investigate QPOs in the burst time series, which provide crucial insights into the underlying radiation mechanisms. Section 4 discusses the implications of our findings, and Section 5 summarizes the main conclusions of this study.
@@
Figure 1: The 57 observations of FAST on FRB 20240114A from January 28 to August 30, 2024, showing the distributions of burst counts (bars in the upper panel), burst rates (red dots in the upper panel), and exposure times (bars in the lower panel). Bar charts with the same color indicate that a consistent short-timescale periodic signal may exist in the corresponding observations (see Section 2).
@@ 2 Searching for Periodic Signals in TOA data
The TOAs in this paper are derived from 57 FAST observations of FRB 20240114A between January 28 and August 30, 2024. For more details on the data processing of this source, including burst searching, dispersion correction, and TOA definition, please refer to the paper by Junshuo Zhang et al. (2025, in preparation). The processed data, including the TOAs and other derived parameters, can be accessed via https://doi.org/10.57760/sciencedb.Fastro.00030. Since the data span several months, PINT (Luo et al., 2021; Susobhanan et al., 2024) was used to convert the TOAs to the solar system barycenter. Figure 1 shows the time distribution of bursts across the 57 observations. We perform periodic signal searches on the TOA data at both short and long timescales.
@@ 2.1.1 Short-Timescale Periodicity Search
Table 1: Candidate periodic signals detected by the TDA and refined using the $Z_{n}^{2}$ statistic. The table lists the candidate number, observation date (MJD), parameters from the TDA detection (frequency $f$, derivative $\dot{f}$, period $P$, derivative $\dot{P}$, significance $\sigma$), and the refined parameters from the $Z_{n}^{2}$ statistic.
@@
Hz ×$\times 10^{-7}$ Hz/s s ×$\times 10^{-7}$ s/s Hz ×$\times 10^{-7}$ Hz/s s ×$\times 10^{-7}$ s/s
1 60401 1.4868 -14.556 0.6726 6.585 4.9 1.4869 -16.301 0.6726 7.374
60403 1.4848 0.000 0.6735 0.000 4.1 1.4854 -8.310 0.6732 3.767
2 60398 1.5759 5.478 0.6346 -2.206 3.4 1.5761 2.304 0.6345 -0.927
60519 1.5739 7.770 0.6354 -3.137 6.0 1.5735 10.384 0.6355 -4.194
3 60417 1.8648 -2.534 0.5363 0.729 3.7 1.8653 -9.304 0.5361 2.674
60517 1.8665 8.259 0.5358 -2.371 3.5 1.8664 8.923 0.5358 -2.561
The TDA has been widely applied in periodicity detection for high-energy gamma-ray pulsars (Atwood et al., 2006; Ziegler et al., 2008; Abdo et. al., 2008; Abdo et al., 2010; Smith et al., 2023). This method is particularly suitable for sparse event signals like FRBs that exhibit only a few burst events per day. The TDA evaluates periodicity by analyzing time differences between events across a sliding time window. The statistical quantity searched by this method is (Atwood et al., 2006)
@@
Our search simultaneously covers both frequency $f$ and its derivative $\dot{f}$, accounting not only for possible spin-down effects of the FRB source but also for potential binary system scenarios, provided that $T_{v}\leq 10P_{\rm b}$, where $P_{\rm b}$ is the orbital period of the FRB source (Ransom et al., 2003). The correction for $\dot{f}$ can be determined through
@@
where $f_{\rm max}$ represents the maximum searched frequency. This parameter space covers plausible physical scenarios for FRB sources. $T_{\rm win}$ balances computational efficiency and frequency resolution: for observations $<$ 1800 s, $T_{\rm win}=T_{v}$; for observations $\geq$ 1800 s, $T_{\rm win}$ is fixed at 1800 s. Previous studies by Wang et al. (2022) suggest that repeating FRBs might originate from magnetars in binary systems with Be stars possessing disk structures. The range of $\dot{f}/f$ explored spans physically plausible regimes for both spin-down and binary motion. In a circular binary orbit, this quantity relates to the orbital period $P_{\rm b}$ and companion mass $M_{\rm c}$ via
@@
Following Niu et al. (2022b), we adopt an upper limit of $M_{c}=100M_{\odot}$ to represent an extreme massive companion. Under this assumption, $\left(\dot{f}/f\right)_{\max}=10^{-6}\ \mathrm{s^{-1}}$, which corresponds to an orbital period of $P_{b}\sim 8.2$ hr. We also consider potential spin-down effects of the FRB source. For magnetars, the magnetic field strength of the surface $B_{s}$ is related to the spin period $P$ and its derivative $\dot{P}$ through $B_{s}=3.2\times 10^{19}\,\mathrm{G}\sqrt{P\dot{P}}$. Assuming a magnetar surface field of $10^{15}\,\mathrm{G}$, the intrinsic frequency evolution yields $\frac{\dot{f}}{f}\sim-9.77\times 10^{-10}f^{2}\ \mathrm{s^{-1}}$. Considering the known spin frequency of magnetars [^1] https://www.physics.mcgill.ca/ pulsar/magnetar/main, the value of $\frac{\dot{f}}{f}$ ranges from $-9.5\times 10^{-9}\ \mathrm{s^{-1}}$ to $-7.13\times 10^{-12}\ \mathrm{s^{-1}}$.
@@
    Dedicated search: The TOAs from each single continuous observation are analyzed separately, covering frequency $f$ and its derivative $\dot{f}$ with $\frac{\dot{f}}{f}$ ranging from $-10^{-6}$ to $10^{-6}\,\mathrm{s}^{-1}$.
@@
    Joint search across days: All TOAs data from multiple observations are searched collectively, with a refined $\frac{\dot{f}}{f}$ range of $-10^{-8}$ to $0\,\mathrm{s}^{-1}$ to focus on long-term spin-down trends.
@@
    Clustered search: Since a single $\dot{f}$ may not fully capture the frequency evolution over long timescales, the TOAs are divided into segments (each $\leq$ 3 days) and analyzed independently. For each segment, $\frac{\dot{f}}{f}$ is constrained to $-10^{-8}$ to $0\,\mathrm{s}^{-1}$, allowing for a more precise characterization of the spin-down behavior within shorter intervals.
@@
where N is the number of TOAs and $\phi_{j}=2\pi(ft_{j}+\frac{1}{2}\dot{f}t_{j}^{2})$. We evaluate the significance of each candidate by computing $Z^{2}_{n}$ over a fine grid around the TDA-derived $(f,\dot{f})$, incorporating the first $n=4$ harmonics to balance sensitivity and computational cost. This two-stage approach, which combines a wide search using the TDA with subsequent $Z^{2}_{n}$ refinement, ensures efficient and reliable detection of weak periodic signals in sparse TOAs.

### v2

# A comprehensive search for Long and Short Periodic Features from an Extremely Active Cycle of FRB 20240114A

@@ Abstract
Possible periodic features in fast radio bursts (FRBs) may provide insights into their astrophysical origins. Using extensive observations from the Five-hundred-meter Aperture Spherical radio Telescope (FAST), we conduct a multi-timescale periodicity search for the exceptionally active repeater FRB 20240114A. Our analysis is based on different datasets for different timescales: for short-timescale periodicity in Time of Arrivals (TOAs), we use 57 observations from January to August 2024; for long-timescale periodicity, we employ an extended TOA dataset comprising 111 observations spanning from January 2024 to October 2025; and for burst time series analysis, we utilize individual burst data from the 57 FAST observations. We identify three candidate short-timescale periodic signals (0.673 s, 0.635 s, and 0.536 s) with significances of $3.2\sigma$$6\sigma$, each detected in two independent observations. On longer timescales, we detect a significant $143.40\pm 7.19$-day periodicity with $5.2\sigma$ significance, establishing FRB 20240114A as a periodic repeater. In burst time series, we find quasi-periodic oscillations in the few hundred Hz range ($3.4\sigma$ and $3.7\sigma$) and periodic burst trains with periods of several to tens of milliseconds ($3\sigma$$3.9\sigma$), though these periodic features appear transient and short-lived. The detection of periodic signals at these different time scales indicates that FRB 20240114A exhibits intriguing periodic self-similar characteristics. Despite the comprehensive dataset, no definitive periodicity linked to the source’s rotation is confirmed, placing stringent constraints on the intrinsic source properties and the modulation mechanisms. All data are available via the Science Data Bank.
keywords:
Fast radio burst, Time series analysis, Magnetars
1 Introduction
Fast radio bursts (FRBs), first discovered in 2007 [lorimer2007bright], are millisecond-duration, highly energetic radio transients of extragalactic origin. They have been detected at cosmological distances and are now recognized as one of the most intriguing phenomena in modern astrophysics [lorimer2007bright, 2013Sci...341...53T, chime/frb2021First, 2022Natur.606..873N]. Since their initial discovery, the field of FRB research has grown rapidly, with extensive observational campaigns leading to the identification of hundreds of these mysterious bursts. These detections have revealed a diverse population, with some FRBs appearing to repeat sporadically while others have only been observed as singular, non-repeating events [2019A&ARv..27....4P, chime/frb2021First, xu2023Blinkverse]. The identification of repeating FRBs, most notably FRB 20121102A [spitler2016repeating], has ruled out the early speculations that all FRBs originate from cataclysmic one-time events. Instead, the existence of repeaters indicates that at least some FRBs are powered by non-catastrophic engines.
A breakthrough in the field of FRB research was marked by the detection of FRB 20200428, which was definitively associated with an X-ray burst from the Galactic magnetar SGR J1935+2154 [2020Natur.587...59B, mereghetti2020INTEGRAL, 2021NatAs...5..378L]. Another targeted multi-wavelength campaign did not detect further coincident radio bursts from this source, suggesting that such FRB-SGR associations are rare and might require extreme physical conditions to produce coherent radio emission [2020Natur.587...63L]. Extensive observations of repeating FRBs have revealed non-Poissonian burst clustering, long-term variations in DM and rotation measure (RM), as well as complex waiting time statistics, highlighting their highly stochastic nature and posing challenges to single rotating compact object models [2021Natur.598..267L, 2023ApJS..269...17H, 2024SciBu..69.1020Z, 2025arXiv250400391L]. Recent polarization and magnetospheric studies suggest that repeating FRBs reside in complex, evolving environments consistent with young magnetars, as indicated by rapid RM variations and strong circular polarization, although observations of FRB 20220912A, which shows high circular polarization but nearly constant RM, suggest that at least some repeaters may not be embedded in strongly magneto-ionic environments [2022Natur.609..685X, 2022Sci...375.1266F, 2023SciA....9F6198Z, 2024ApJ...974..296F, 2025SCPMA..6889511F]. A successful search for pulsar-like, spin-modulated burst-to-burst temporal periodicities in FRB sources would provide a convincing observational evidence that at least some FRBs originate from compact stars.
Some FRB sources have demonstrated distinct long-timescale periodic activity. Notable examples include FRB 20180916B, which follows a 16-days cycle of activity [2020Natur.582..351C, 2021Natur.596..505P, 2023ApJ...956...23S], and FRB 20121102A, which has shown signs of a much longer periodicity of approximately 160 days, alongside a shorter periodicity of around 4.6 days [2020MNRAS.495.3551R, 2021MNRAS.500..448C, 2024ApJ...969...23L, 2025A&A...693A..40B]. The recently discovered FRB 20240209A, located at the outskirts of a quiescent elliptical galaxy (z = 0.1384), exhibits a periodic activity cycle of about 126 days, further adding to the rare class of periodically repeating FRBs [2025ApJ...983L..15P]. Furthermore, FRB 20220529 shows a potential periodicity of approximately 200 days in its RM, which is consistent with a binary origin, as suggested by the significant RM variation and its rapid recovery [2025arXiv250510463L]. Similarly, FRB 20201124A exhibits a periodicity of 26 days in its RM, further supporting the idea that periodic RM variations may be a common feature of actively repeating FRBs in binary systems [2025arXiv250506006X]. The origin of periodic activity in repeating FRBs remains debated, with proposed explanations falling into several categories. Rotation-powered models suggest the periodicity reflects the spin of isolated magnetars [2021ApJ...917....2X, 2024ApJ...967L..44L]. Binary models attribute the modulation to orbital dynamics in neutron star systems, including Be/X-ray binaries [2020ApJ...893L..26I, 2020ApJ...893L..39L, 2020MNRAS.498L...1Z, 2021ApJ...918L...5L, 2022MNRAS.515.4217B]. Precession effects have also been proposed as a possible cause of the observed periodicity [2020ApJ...893L..31Y, 2020RAA....20..142T, 2020ApJ...895L..30L, 2020ApJ...892L..15Z, 2020MNRAS.494L..64K, 2020PASJ...72L...8C, 2021ApJ...921..147C, 2021ApJ...917...13S, 2022A&A...658A.163W]. Alternative explanations involve asteroid belt interactions [2020ApJ...895L...1D, 2021MNRAS.508.2079V, 2024ApJ...974..215D] or the repetitive collapse of the crust of an accreting strange quark star [2021Innov...200152G]. The diversity of models highlights the need for further observational constraints, as many fundamental questions about the physical origins and emission mechanisms of FRBs remain open [2023RvMP...95c5005Z].
Short-timescale periodicity in FRBs provides critical constraints on their emission mechanisms and progenitors. Some bursts have shown millisecond-scale quasi-periodic structures, e.g. the 0.411 ms spacing in FRB 20201020A [2023A&A...678A.149P], although its significance remains marginal. These sub-second periodicities challenge the spin explanations due to extreme rotational requirements, favoring magnetospheric processes instead. Additionally, the detection of ultra-narrow spike components in FRB 20230607A, down to 0.3 ms, further supports a magnetospheric origin for at least some repeating FRBs [2025arXiv250411173Z]. On the other hand, other studies have found no significant periodicity in repeating FRBs like FRB 20201124A, ruling out isolated magnetars or binaries with certain parameters [2022RAA....22l4004N]. Recent claims of a $\sim$1.7 s period in FRB 20201124A [2025arXiv250312013D], if confirmed, could indicate a pulsar origin, while low-significance periodicity candidates such as the 7.267 ms signal in FRB 20230708A [2025MNRAS.536.3220D] highlight the challenges in distinguishing true periodicity from burst microstructure. These controversial results highlight the need for higher-cadence and higher time resolution observations to clarify the nature of short-timescale FRB variability.
@@ 1 Introduction
This paper is organized as follows. In Section 2, we search for periodic signals in the time-of-arrival (TOA) data of FRB 20240114A to explore possible links to magnetar rotation, binary orbits, and precession scenarios. In Section 3, we mainly investigate QPOs in the burst time series, which provide crucial insights into the underlying radiation mechanisms. Section 4 discusses the implications of our findings, and Section 5 summarizes the main conclusions of this study.
@@
Figure 1: The 57 FAST observations of FRB 20240114A from January 28 to August 29, 2024 (UT), show the distribution of waiting times between bursts (top panel), burst counts (bars in the middle panel), burst rates (red dots in the middle panel), cumulative burst counts (dashed line in the middle panel), and exposure times (bars in the bottom panel). In the top panel, vertical lines of different colors represent the period values of three candidates identified in the TDA short-timescale search. The bar charts with the same colors in the middle panel indicate that there are candidate periodic signals with similar parameters in the corresponding observations (see Section 2).
@@ 2 Searching for Periodic Signals in TOA data
The primary TOAs analyzed in this paper, which are used for most analyses including the short-timescale periodicity search, are derived from 57 FAST observations of FRB 20240114A between January 28, 2024, and August 29, 2024 (UT). Figure 1 shows the time distribution of bursts across these 57 observations, as well as the distribution of waiting times between consecutive TOAs within this dataset.
To enhance the sensitivity of the long-timescale periodicity search to multi-month activity cycles, we specially extend the dataset to include TOAs from observations up to October 1, 2025 (UT). While the short-timescale analyses rely on the original dataset for high-precision single-pulse measurements suitable for detecting minute-to-hour periodicities, the extended dataset provides reliable TOAs and burst rates appropriate for detecting long-term activity cycles. For more details on the data processing of this source, including burst searching, dispersion correction, and TOA definition, please refer to [zhang2025a]. The TOAs were converted to the solar system barycenter using PINT [2021ApJ...911...45L, 2024ApJ...971..150S].
@@ 2.1.1 Short-Timescale Periodicity Search
Table 1: Candidate periodic signals detected by the TDA method and refined using the $Z_{n}^{2}$ statistic. The table lists the candidate number, observation date (MJD), parameters from the TDA detection (frequency $f$, derivative $\dot{f}$, period $P$, derivative $\dot{P}$, significance $\sigma$), and the refined parameters from the $Z_{n}^{2}$ statistic.
@@
Hz $\times 10^{-7}$ Hz/s s $\times 10^{-7}$ s/s Hz $\times 10^{-7}$ Hz/s s $\times 10^{-7}$ s/s
1 60401 1.48677(58) -14.322(78) 0.67260(26) 6.479(36) 4.9 1.4867871(30) -14.319070(29) 0.6725912(13) 6.477646(29)
60403 1.48482(58) -1.398(78) 0.67348(26) 0.634(35) 4.2 1.4849567(30) -1.3992047(28) 0.6734203(13) 0.6345322(28)
2 60398 1.57588(59) 5.138(85) 0.63457(24) -2.069(34) 3.2 1.5761418(32) 5.133988(10) 0.6344607(13) -2.0666374(93)
60519 1.57392(59) 7.686(85) 0.63536(24) -3.103(34) 6.0 1.5737813(32) 7.692812(15) 0.6354123(13) -3.105964(14)
3 60417 1.86477(58) -2.729(98) 0.53626(17) 0.785(28) 3.6 1.8648264(37) -2.7264407(55) 0.5362429(11) 0.7840057(35)
60517 1.86652(58) 8.549(97) 0.53576(17) -2.454(28) 3.5 1.8665689(37) 8.557654(17) 0.5357423(11) -2.456216(11)
The TDA has been widely applied in periodicity detection for high-energy gamma-ray pulsars [2006ApJ...652L..49A, 2008ApJ...680..620Z, 2008Sci...322.1218A, 2010ApJS..187..460A, 2023ApJ...958..191S]. In particular, Atwood et al. ([2006ApJ...652L..49A]) demonstrated that TDA can achieve comparable sensitivity to FFT-based searches while being over an order of magnitude faster for sparse TOA data, making it especially suitable for FRBs, whose bursts are typically sparsely and irregularly distributed over the observation period. The TDA evaluates periodicity by analyzing time differences between events across a sliding time window. The statistical quantity searched by this method is [2006ApJ...652L..49A]
@@
Our search simultaneously covers both frequency $f$ and its derivative $\dot{f}$, accounting not only for possible spin-down effects of the FRB source but also for potential binary system scenarios, provided that $T_{v}\leq 10P_{\rm b}$, where $P_{\rm b}$ is the orbital period of the FRB source [2003ApJ...589..911R]. The correction for $\dot{f}$ can be determined through
@@
where $f_{\rm max}$ represents the maximum searched frequency. This parameter space covers plausible physical scenarios for FRB sources. $T_{\rm win}$ balances computational efficiency and frequency resolution: for observations << 1800 s, $T_{\rm win}=T_{v}$; for observations $\geq$ 1800 s, $T_{\rm win}$ is fixed at 1800 s. Previous studies by the reference [2022NatCo..13.4382W] suggest that repeating FRBs might originate from magnetars in binary systems with Be stars possessing disk structures. The range of $\dot{f}/f$ explored spans physically plausible regimes for both spin-down and binary motion. In a circular binary orbit, this quantity relates to the orbital period $P_{\rm b}$ and companion mass $M_{\rm c}$ via
@@
Following the reference [2022RAA....22l4004N], we adopt an upper limit of $M_{c}=100M_{\odot}$ to represent an extreme massive companion. Under this assumption, $\left(\dot{f}/f\right)_{\max}=10^{-6}\ \mathrm{s^{-1}}$, which corresponds to an orbital period of $P_{b}\sim 8.2$ hr. We also consider potential spin-down effects of the FRB source. For magnetars, the magnetic field strength of the surface $B_{s}$ is related to the spin period $P$ and its derivative $\dot{P}$ through $B_{s}=3.2\times 10^{19}\,\mathrm{G}\sqrt{P\dot{P}}$. Assuming a magnetar surface field of $10^{15}\,\mathrm{G}$, the intrinsic frequency evolution yields $\frac{\dot{f}}{f}\sim-9.77\times 10^{-10}f^{2}\ \mathrm{s^{-1}}$. Considering the known spin frequency of magnetars [^7] https://www.physics.mcgill.ca/ pulsar/magnetar/main, the value of $\frac{\dot{f}}{f}$ ranges from $9.5\times 10^{-9}\ \mathrm{s^{-1}}$ to $7.13\times 10^{-12}\ \mathrm{s^{-1}}$.
@@
    Dedicated search: The TOAs from each single continuous observation are analyzed separately, covering frequency $f$ and its derivative $\dot{f}$ with $\frac{\dot{f}}{f}$ ranging from $10^{-6}$ to $10^{-6}\,\mathrm{s}^{-1}$.
@@
    Joint search across days: All TOAs data from multiple observations are searched collectively, with a refined $\frac{\dot{f}}{f}$ range of $10^{-8}$ to $0\,\mathrm{s}^{-1}$ to focus on long-term spin-down trends.
@@
    Clustered search: Since a single $\dot{f}$ may not fully capture the frequency evolution over long timescales, the TOAs are divided into segments (each $\leq$ 3 days) and analyzed independently. For each segment, $\frac{\dot{f}}{f}$ is constrained to $10^{-8}$ to $0\,\mathrm{s}^{-1}$, allowing for a more precise characterization of the spin-down behavior within shorter intervals.
@@
where N is the number of TOAs and $\phi_{j}=2\pi(ft_{j}+\frac{1}{2}\dot{f}t_{j}^{2})$. We evaluate the significance of each candidate by computing $Z^{2}_{n}$ over a fine grid around the TDA-derived $(f,\dot{f})$, incorporating the first $n=4$ harmonics to balance sensitivity and computational cost. This two-stage approach, which combines a wide search using the TDA with subsequent $Z^{2}_{n}$ refinement, ensures efficient and reliable detection of weak periodic signals in sparse TOAs. A demonstration of this TDA and $Z_{n}^{2}$ pipeline on simulated TOA data is provided in A.
Figure 2: Folding results of the short-timescale periodicity candidates. Each row corresponds to one candidate group, with the left and right panels showing similar periods observed on different observation dates (see Table 1). Top of each panel indicates the candidate period $P$ in seconds. For each candidate, the top panel displays the pulse profile folded with the corresponding period after correcting for the period derivative, and the bottom panel shows the distribution of individual TOAs in phase.